calculate the grams of NaCl in a sample containing 1.2x10^24
1
Expert's answer
2020-12-08T05:34:12-0500

1 mole of NaCl = 6.02×10²³ molecules

58.5g of NaCl = 6.02×10²³ molecules

xg of NaCl = 1.2×10²⁴ molecules


x = "\dfrac{1.2×10²⁴×58.5}{6.02×10²³}"


= 2×58.5g = 117g


Therefore, 1.2×10²⁴ molecules of NaCl has a mass of 117g.